An object moves along the axis according to the equation x(t)= (3.002 - 2.00+ 3.00) m, where is in seconds. Determine

(a) the instantaneous velocity at = 2.00 s and = 3.00 s, 

(b) the average acceleration between = 2.00 s and = 3.00 s,

(c) the instantaneous acceleration at = 3.00 s



1
Expert's answer
2021-09-17T17:34:59-0400

a)

"v=\frac{dx}{dt}=6t-2,"

"v(t=2)=10~\frac ms,"

"v(t=3)=16~\frac ms,"

b)

"a=\frac{v_3-v_2}{t_3-t_2}=\frac{16-10}{3-2}=6~\frac m{s^2},"

c)

"a=\frac{dv}{dt}=6~\frac m{s^2}."